OpenStudy (notamathgenius):

write an equation in slope-intercept form of the line that passes through the given points (3, -2), (6,1) @ajprincess Same way?

OpenStudy (ajprincess):

have u found the slope using that formula?

OpenStudy (notamathgenius):

No c: i know nothing about slope intercept I dislike it a lot -_-

OpenStudy (ajprincess):

slope is found using the formula \[slope~m=\frac{y_2-y_1}{x_2-x_1}\] \[m=\frac{1-(-2)}{6-3}=?\]

OpenStudy (notamathgenius):

i got -1.666666667

OpenStudy (ajprincess):

nw can u find m?

OpenStudy (ajprincess):

@notamathgenius

OpenStudy (notamathgenius):

hmmm

OpenStudy (ajprincess):

sure?

OpenStudy (notamathgenius):

Yea I am lost, i keep looking back at the question -_- grr i hate math

OpenStudy (ajprincess):

wht is 1-(-2)=?

OpenStudy (notamathgenius):

3

OpenStudy (ajprincess):

good:) nw what is 6-3=?

OpenStudy (notamathgenius):

3

OpenStudy (ajprincess):

good:) 3/3=?

OpenStudy (notamathgenius):

1

OpenStudy (ajprincess):

good:) so what is m=?

OpenStudy (notamathgenius):

1 :D

OpenStudy (ajprincess):

yup. good:)

OpenStudy (notamathgenius):

Oh boy math is fun -_- not

OpenStudy (ajprincess):

nw let us consider a point (x.y) on the same line. then \[m=\frac{y-y_1}{x-x_1}\] \[1=\frac{y-(-2)}{x-3}\] \[1(x-3)=y+2\] Slope-intercept equation is \\[y=1(x-3)-2\]

OpenStudy (notamathgenius):

I see now :) thanks a bunch

OpenStudy (ajprincess):

welcome:) remove the parentheses and simplify:)

OpenStudy (notamathgenius):

WIll do :D

OpenStudy (ajprincess):

good:)